# MissionHD: Hyperdimensional Refinement of Distribution-Deficient Reasoning Graphs for Video Anomaly Detection

### v3

@@ 5.2 Quantitative Evaluation of MissionHD
We quantitatively evaluated our method against existing graph refinement approaches by applying them to LLM-generated MSGs. The baselines include four strong methods: direct optimization of the adjacency matrix, two graph neural networks (GCN kipf2017semi and GAT velivckovic2017graph ), and the state-of-the-art energy-based refinement method ECL-GSR zeng2025graph .
@@ Appendix C Empirical Analysis of Task-Driven Graph Refinement
Robbery Pruned (General / Abstract)
theft, danger, illegal, crime, fear
Retained (Specific Objects / Actions)
weapon, mask, getaway car, forced entry,
cash register, demanding money
Vandalism Pruned (General / Abstract)
damage, destruction, malicious, illegal act
Retained (Specific Objects / Actions)
graffiti, spray paint, breaking windows,
smashing, keying car
@@
Top-Scored Edges
(Visual Cue $\rightarrow$ Implication)
3.51% 3.45% 3.38%
weapon $\rightarrow$ armed robbery
mask $\rightarrow$ identity concealed
forced entry $\rightarrow$ property crime
Bottom-Scored Edges
(Abstract $\rightarrow$ Abstract)
1.88% 1.81% 1.75%
danger $\rightarrow$ high stress
illegal $\rightarrow$ criminal intent
theft $\rightarrow$ financial loss

### v4

@@ 5.2 Quantitative Evaluation of MissionHD
We quantitatively evaluated our method against existing graph refinement approaches by applying them to LLM-generated MSGs. The baselines include four strong methods: direct optimization of the adjacency matrix, two graph neural networks (GCN Kipf and Welling (2017) and GAT Veličković et al. (2017)), and the state-of-the-art energy-based refinement method ECL-GSR Zeng et al. (2025b).
@@ Appendix C Empirical Analysis of Task-Driven Graph Refinement
Robbery Pruned (General / Abstract) theft, danger, illegal, crime, fear
Retained (Specific Objects / Actions) weapon, mask, getaway car, forced entry, cash register, demanding money
Vandalism Pruned (General / Abstract) damage, destruction, malicious, illegal act
Retained (Specific Objects / Actions) graffiti, spray paint, breaking windows, smashing, keying car
@@
Top-Scored Edges (Visual Cue $\rightarrow$ Implication) 3.51% 3.45% 3.38% weapon $\rightarrow$ armed robbery mask $\rightarrow$ identity concealed forced entry $\rightarrow$ property crime
Bottom-Scored Edges (Abstract $\rightarrow$ Abstract) 1.88% 1.81% 1.75% danger $\rightarrow$ high stress illegal $\rightarrow$ criminal intent theft $\rightarrow$ financial loss